OpenStudy (lolipop_drum312):

What is the coefficient in the expression 2x/7+2 ? my answer 2/7 is this right Enter your answer in the box.

OpenStudy (lolipop_drum312):

@3mar

OpenStudy (3mar):

Yes.

OpenStudy (3mar):

\[\frac{ 2x }{ 7 }+2\]??

OpenStudy (lolipop_drum312):

yes

OpenStudy (lolipop_drum312):

ok next one

OpenStudy (lolipop_drum312):

How many days equal 5616 minutes? my answer 3.9 Enter your answer in the box. 5616 min = day

OpenStudy (3mar):

my answer 2/7 is this right Yes, this is right.

OpenStudy (lolipop_drum312):

ok next and idk yet

OpenStudy (3mar):

Ok

OpenStudy (lolipop_drum312):

How many minutes equal 5.4 days? Enter your answer in the box. 5.4 days = 7776 min < my answer

OpenStudy (civicsiscool44):

Yes you are right

OpenStudy (lolipop_drum312):

ok next

OpenStudy (civicsiscool44):

medal me please

OpenStudy (lolipop_drum312):

um cn you paste me there

OpenStudy (lolipop_drum312):

A turtle swims at a speed of 15,840 yd per h. A girl swims at a speed of 138 cm per s. 1 dekameter (dam) = 11 yd 1 dam = 1000 cm How much faster does the turtle swim than the girl in dekameters per minute? Enter your answer in the box. 15.9 dammin < my answer

OpenStudy (lolipop_drum312):

dam/min*

563blackghost (563blackghost):

I agree with @3mar you are right :)

OpenStudy (lolipop_drum312):

and 15,9 is my answer is it right

OpenStudy (3mar):

to move from minute to year: \[\frac{ 7776 \min }{ hr }=129.6 hr\] then \[\frac{ 129.6hr }{ 24 }=5.4 days\]

OpenStudy (lolipop_drum312):

ok so it is 7776

OpenStudy (lolipop_drum312):

now did i get the toher one right

OpenStudy (3mar):

Can we start a post for one question?

OpenStudy (lolipop_drum312):

a new post you mean

OpenStudy (3mar):

yes

OpenStudy (lolipop_drum312):

ok

OpenStudy (3mar):

As it take too long to reach from top to down

OpenStudy (3mar):

Thank you for the medals!